OpenStudy (anonymous):

The length of a rectangle. Is five feet longer than its width. If the perimeter of the rectangle is 42 feet find its area.

OpenStudy (pratyush5):

let width = x metres therefore length = x+5 metres No we know perimeter of a rectangle = 2 (l+b) therefore, 2 ( (x+5)+(x) ) = 42 Solve for x and find length and breadth Then calculate area :)

OpenStudy (anonymous):

thanks

OpenStudy (pratyush5):

You're welcome :)

OpenStudy (pratyush5):

So what area do you get ?